OpenStudy (anonymous):

Can someone please help me (will give medal and fan) Given the following system of equations, identify the type of system. x + y = 6 y = 3 - x consistent inconsistent equivalent

OpenStudy (anonymous):

Solve the following system of equations graphically. x + y - 4 = 0 x - y = 0 The solution lies in quadrant _____. I II III IV

OpenStudy (anonymous):

Given the following system of equations, identify the type of system. x + y = 4 x - y = 6 consistent inconsistent equivalent

OpenStudy (anonymous):

Solve the following system of equations graphically. y - 4 = 0 2x - y - 2 = 0 What is the solution set? {(-3, -4)} {(-3, 4)} {(3, 4)}

OpenStudy (anonymous):

Solve the following system of equations graphically. x = -5 y = -6 What is the solution set? {(-6, -5)} {(-5, -6)} Ø

OpenStudy (anonymous):

Solve the following system of equations graphically. y = 2x - 3 x + y = 3 What is the solution set? {(2, 1)} {(1, 2)} {(-1, -2)} {(-2, -1)}